### v1

# CausalRAG: Integrating Causal Graphs into Retrieval-Augmented Generation

Abstract
Large language models (LLMs) have revolutionized natural language processing (NLP), particularly through Retrieval-Augmented Generation (RAG), which enhances LLM capabilities by integrating external knowledge. However, traditional RAG systems face critical limitations, including disrupted contextual integrity due to text chunking, and over-reliance on semantic similarity for retrieval. To address these issues, we propose CausalRAG, a novel framework that incorporates causal graphs into the retrieval process. By constructing and tracing causal relationships, CausalRAG preserves contextual continuity and improves retrieval precision, leading to more accurate and interpretable responses. We evaluate CausalRAG against regular RAG and graph-based RAG approaches, demonstrating its superiority across several metrics. Our findings suggest that grounding retrieval in causal reasoning provides a promising approach to knowledge-intensive tasks.
CausalRAG: Integrating Causal Graphs into Retrieval-Augmented Generation
Nengbo Wang1,2, Xiaotian Han1, Jagdip Singh2, Jing Ma1, Vipin Chaudhary1 1Department of Computer and Data Sciences, Case Western Reserve University 2Department of Design and Innovation, Case Western Reserve University nengbo.wang@case.edu
1 Introduction
The rapid advancements in large language models (LLMs) have revolutionized the field of natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024). However, their reliance on pre-trained knowledge limits their ability to integrate and reason over dynamically updated external information, particularly in knowledge intensive domains such as academic research. Retrieval-Augmented Generation (RAG) has emerged as a promising framework to address this limitation Lewis et al. (2021), combining retrieval mechanisms with generative capabilities to enhance contextual understanding and response quality.
Recent research has focused on improving RAG along two primary directions: 1) enhancing retrieval efficiency and integration mechanisms by designing more adaptive and dynamic retrieval frameworks Gan et al. (2024); Ravuru et al. (2024); Zhang et al. (2024a); 2) improving the representation of external knowledge to facilitate retrieval and reasoning, with graph-based RAGs being a dominant approach Edge et al. (2024); Guo et al. (2024); Potts (2024). Despite these advancements, existing RAG architectures still face critical limitations that impact retrieval quality and response accuracy, primarily due to three key issues: 1) disruption of contextual integrity caused by the text chunking design; 2) reliance on semantic similarity rather than causal relevance for retrieval; and 3) a lack of accuracy in selecting truly relevant documents.
Through a combination of theoretical analysis and empirical evaluation, we rethink the limitations of current RAG systems by introducing a novel perspective based on context recall and precision metrics. Our findings reveal that both regular and graph-based RAGs struggle not only to retrieve truly grounded context but also to accurately discern the relationship between retrieved content and the user query. We identify this fundamental issue as one primary reason why LLMs in RAG frameworks often generate seemingly relevant yet shallow responses that lack essential details.
To address these gaps, we introduce CausalRAG, a novel RAG framework that integrates causal graphs to enhance retrieval accuracy and reasoning performance. Unlike regular and graph-based RAGs, CausalRAG constructs a causal graph from uploaded documents, preserving contextual relationships while capturing cause-effect dependencies. By ensuring that retrieved documents are both relevant and causally grounded, CausalRAG enables the generation of more contextually rich and causally detailed responses. This approach not only improves retrieval effectiveness but also mitigates hallucinations and enhances answer faithfulness.
We evaluate CausalRAG on datasets from diverse domains and across varying context lengths, comparing its performance with regular RAG and GraphRAG, a graph-based RAG framework from Microsoft Edge et al. (2024). Our experiments assess performance across three key metrics: answer faithfulness, context recall, and context precision. Results demonstrate that CausalRAG achieves superior performance across different contexts. Additionally, we conduct a case study and a parameter analysis to further examine our framework, analyzing and providing insights that contribute to ongoing research in RAG. The contributions of this work are threefold:
  • We systematically identify the inherent limitations of RAG’s retrieval process through analytical and experimental study. More importantly, we uncover one major reason why LLMs in RAG tend to generate superficial, generalized answers that lack the grounded details expected by users.
  • We propose CausalRAG, a framework that enhances both retrieval and generation quality by incorporating causality into the RAG, effectively addressing these limitations.
  • Our work further mitigates hallucination issues and significantly improves the interpretability of AI systems. We summarize key findings and insights in both retrieval and generation process, contributing to research in RAG.
2 Related Work
2.1 Retrieval-Augmented Generation
RAG enhances LLMs’ ability to handle knowledge-intensive tasks by integrating external knowledge retrieval Lewis et al. (2021). Existing research primarily advances RAG along two key dimensions: 1) improving retrieval efficiency and integration mechanisms; 2) enhancing the representation of external knowledge to facilitate reasoning and interpretability.
Optimizing retrieval flow and interaction. The first stream focuses on improving the interaction flow within the RAG system to enhance output quality. Approaches have introduced pre-retrieval, retrieval, and post-retrieval refinements to mitigate redundancy and computational overhead Wang et al. (2024). Modular RAG architectures further advance this by enabling iterative retrieval-generation cycles, allowing dynamic interactions between retrieval and content creation. For example, CAiRE-COVID Su et al. (2020) demonstrated the effectiveness of iterative retrieval in multi-document summarization, while some work Feng et al. (2023) extended this approach to multi-hop question answering. Recent innovations include METRAG Gan et al. (2024), which integrates LLM supervision to generate utility-driven retrieval processes, and RAFT Zhang et al. (2024a), which trains models to disregard distractor documents while improving citation accuracy through chain-of-thought reasoning.
Structuring external knowledge for efficiency. The second direction explores improved methods for structuring external knowledge to achieve better retrieval efficiency. For example, GraphRAG Edge et al. (2024), proposed by Microsoft, treats external knowledge as interconnected nodes, capturing causal and thematic relationships to enhance retrieval depth and reasoning. LightRAG introduced a dual-level retrieval mechanism for incremental knowledge updates Guo et al. (2024). More recently, Lazy GraphRAG has been developed as a continuous version of GraphRAG, deferring computationally expensive operations until query time and leveraging lightweight indexing techniques for efficiency Potts (2024). Despite these advancements, ensuring the quality and relevance of retrieved documents remains a significant concern in current RAG systems, as it directly impacts the coherence of the final response Gupta et al. (2024).
2.2 Causal Graphs and RAG
The combination of causal graphs and RAG has emerged as a promising approach to enhancing knowledge retrieval and reasoning. As causality provides a structural understanding of dependencies within data, it enables more interpretable and reliable AI outputs Ma (2024). Existing research in this domain primarily advances causal discovery with RAG and LLMs. For instance, some work proposed an LLM-assisted breadth-first search (BFS) method for full causal graph discovery, significantly reducing time complexity Jiralerspong et al. (2024). Additionally, some further introduced a correlation-to-causation inference (Corr2Cause) task to evaluate LLMs’ ability to infer causation from correlation, revealing their limitations in generalization across different datasets (Jin et al., 2024).
Despite these advancements, most studies focus on utilizing RAG or LLMs for causal discovery or causal effect estimation Ma (2024); Kıcıman et al. (2024), whereas the direct integration of causal graphs into RAG architectures remains largely unexplored. Our work aims to be a pioneer in this direction. A few existing studies have touched upon this concept but differ in scope. One approach integrates causal graphs within the LLM architecture itself, structuring the transformer’s internal token processing using causality rather than enhancing RAG retrieval Chen et al. (2024b). Another employs causal graphs in RAG systems but focuses on the pre-retrieval stage and largely reduces the core process into a single embedding model without deeper exploration Samarajeewa et al. (2024). GraphRAG is a well-regarded and influential work in this area, as it introduces a graph-based structure into the RAG system, leveraging graph community detection and summarization techniques for retrieval Edge et al. (2024). While it does not incorporate causality, it has significantly improved RAG performance. Therefore, we adopt GraphRAG as the baseline for our work.
In the following sections, we first analyze the nature of regular RAG and grah-based RAGs from a novel perspective, identifying their inherent limitations. We then introduce a causal graph structure to address these gaps within the RAG system and present our framework—CausalRAG.
Figure 1: Analytical and experimental studies reveal limitations in regular RAG and GraphRAG. (a) identifies three key retrieval and generation issues in regular RAG; (b) evaluates RAG via context precision and recall, showing regular RAG excels in recall but lacks precision. GraphRAG improves precision but trades off some recall.
3 Why Regular RAG Fails in Providing Accurate Responses
In this section, through both analytical and experimental investigations, we identify three fundamental limitations of regular RAG and rethink its design by examining its three core elements—user query, retrieved context, and response—through a novel perspective based on precision and recall.
3.1 Limitations of regular RAG
The first limitation arises from RAG’s common practice of chunking texts into minimal units (as illustrated in Figure 1a). This process disrupts the natural linguistic and logical connections in the original text. These connections are crucial for maintaining contextual integrity, and if they are lost, an alternative mechanism must be implemented to restore them.
The second limitation lies in the semantic search process. RAG typically retrieves the semantically closest documents from a vector database based on query similarity. However, in many cases, critical information necessary for answering a query is not semantically similar but rather causally relevant. A classic example is the relationship between diapers and beer—while they are not semantically related, they may exhibit a causal connection in real worlds. This limitation suggests that RAG’s reliance on semantic similarity may lead to the retrieval of contextually irrelevant but superficially related information.
The third limitation is that even when RAG retrieves a relevant context, this does not necessarily guarantee an accurate response. To formalize this issue, we used two key metrics: context recall and context precision, defined as follows:
Context $\text{Context Recall}=\frac{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}{|R|}$ (1)
Here $R$ is the reference set of all relevant reference documents. $C_{i}$ is the $i^{th}$ retrieved document. $\mathbb{I}(C_{i}\in R)$ is an indicator function that returns 1 if $C_{i}$ belongs to the reference set $R$, otherwise 0. It should always return 1 if no hallucination occurs in the LLM.
Context $\text{Context Precision}=\frac{\sum_{i=1}^{N}\mathbb{I_{Q}}(C_{i}\in R)}{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}$ (2)
Here $\mathbb{I_{Q}}(C_{i}\in R)$ is an indicator function that returns 1 if the context retrieved is causally related to user’s query, otherwise 0.
Recall-Precision Perspective. Context recall measures how much of the correct contextual information can be retrieved from the user’s uploaded documents given a query. While RAG primarily focuses on retrieving related documents, adjusting retrieval parameters to increase the number of retrieved documents can generally lead to higher recall in reality. However, context precision presents a challenge. It measures the proportion of retrieved documents that are actually correct given the user query. As discussed earlier, RAG’s reliance on semantic similarity rather than causal relevance often leads to the retrieval of superficially similar but logically irrelevant content. In summary, while RAG can recall numerous answers from reference materials, the proportion of correct responses remains low, ultimately reducing its precision. This recall-precision perspective provides a new lens to see limitations of regular RAG.
3.2 Rethinking Graph-based RAGs
Applying this perspective, we can better understand why Graph-based RAGs serve as improved variants of RAG. By summarizing and ranking the importance of graph communities before retrieval, they largely enhance the quality of retrieved context, thereby improving context precision. However, it only partially addresses the identified limitations, as its summarization process does not entirely filter out irrelevant information. More critically, its reliance on community-based summarization for retrieval may adversely impact recall. Based on these analytical insights, we further conducted an experimental study to validate our analysis.
Experimental study. As shown in Figure 1(b), we conducted an experimental study to empirically verify our analysis. Specifically, we employed an LLM to evaluate the performance of regular RAG and GraphRAG using the RAG evaluation framework Ragas Es et al. (2023). It takes as input the three key elements—user query, retrieved context, and reference—along with predefined metric definitions, and prompts the LLM to assign numerical ratings. This LLM-based evaluation has been widely used in recent RAG research and has also been adopted in RAG evaluations Samarajeewa et al. (2024); Edge et al. (2024).
The results indicate that both global and local versions of GraphRAG achieve higher context precision compared to regular RAG. However, its recall performance remains unsatisfactory, even slightly lower than that of RAG, due to its graph community-based retrieval process.
By combining our precision-recall-based analytical analysis with empirical findings, we clearly outline the inherent limitations of both RAG and its graph-based extensions. In the next section, we introduce our proposed framework, CausalRAG, designed to address these issues.
Figure 2: Overview of CausalRAG’s architecture. Documents are indexed as graphs, and queries retrieve causally related nodes. A causal summary is generated and combined with the query to ensure grounded responses.
4 Methodology
In this section, we introduce our novel framework—CausalRAG—which integrates RAG with causality to overcome the limitations of existing RAG systems. Overall, CausalRAG utilizes a graph-based approach to represent uploaded documents, enhancing robustness against discontinuities caused by text chunking (as shown in Figure 2). More importantly, by expanding and tracing nodes in the graph through causal relationships, CausalRAG is able to retrieve a more diverse set of contextual information while maintaining causal groundedness. This enables CausalRAG to achieve both high recall and strong precision. We now discuss each step in detail.
4.1 Indexing
At the outset, once the system receives the user’s uploaded documents and query, we first index these inputs into our vector database. For the uploaded documents, we employ a text-based graph construction method to transform the text into a structured graph before storing the nodes and edges in the vector database. Specifically, we leverage an LLM to construct the graph following the approach by LangChain Chase (2022), where the LLM scans the text to identify graph nodes and determine relationships between them. Although LLM-based graph construction has shown strong performance Chen et al. (2024a) and is widely adopted in graph-based RAG research , we further validate our constructed graph using expert knowledge, which we elaborate on in the case study. After constructing and indexing the base graph, we embed the user query into the vector database, preparing it for subsequent search and matching. Notably, this indexing process occurs independently of query time, allowing for efficient retrieval during inference.
4.2 Discovering and Estimating Causal Paths
This is the first step during query time. We begin by matching the user query to nodes in the graph based on their embedding distance. The $k$ nodes with the smallest distances are selected, representing the most relevant information directly related to the query. Notably, $k$ is a tunable parameter—higher values retrieve more relevant information at the cost of increased computational complexity.
After selecting the initial $k$ nodes, we expand the search along the base graph’s edges by a step size of $s$, thereby broadening the retrieved context. This step is crucial as it preserves causal and relational connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall. The parameter $s$ controls the depth of expansion, where higher values lead to more diverse information retrieval.
Once the relevant nodes and edges are collected, we employ an LLM to identify and estimate causal paths within them, constructing a refined causal graph. As discussed, LLMs have demonstrated efficiency in discerning and analyzing causal relationships Zhang et al. (2024b); Zhou et al. (2024). This step ensures that CausalRAG prioritizes causally relevant information, improving precision.
Furthermore, the derived causal graph serves two key purposes: 1) It preserves causally relevant information that traditional retrieval methods struggle to capture. More importantly, by adjusting the parameter $s$, this approach can capture long-range causal relationships within the text, particularly when the text is lengthy; 2) It filters out semantically related but causally irrelevant information. Without this filtering, responses may contain unnecessary or even hallucinated content, compromising answer faithfulness.
4.3 Retrieving Context Causally
After constructing the causal graph, we summarize the retrieved information and generate a causal summary. Notably, the input at this stage consists of information that is not only highly relevant but also causally grounded in the user’s query, ensuring greater validity. This approach contrasts with traditional retrieval methods, which often rely purely on semantic similarity and may retrieve contextually related yet causally irrelevant information.
The causal summary is derived by tracing key causal paths within the graph, prioritizing nodes and relationships that contribute directly to answering the query. This ensures that the retrieved information maintains logical coherence and factual consistency while filtering out spurious or weakly related context. Additionally, by leveraging causal dependencies, our method reduces the risk of retrieving semantically similar but misleading evidence.
Once the causal summary is generated, it is combined with the user query to construct a refined prompt for CausalRAG. This structured final input allows RAG to focus on reasoning through causal relationships rather than merely aggregating loosely related text spans.
5 Experiment
To evaluate the effectiveness of CausalRAG, we conduct a series of experiments comparing it with regular RAG and GraphRAG baselines. Our evaluation spans multiple datasets, retrieval settings, and performance metrics, ensuring a comprehensive analysis of retrieval quality and answer faithfulness. We systematically explore different parameter settings and retrieval strategies to assess their impact on model performance.
5.1 Experimental Setup
Parameters & Baselines. We set the default parameters for CausalRAG as $k=s=3$ and use the same $k$ value for GraphRAG’s community-based retrieval and regular RAG’s document retrieval to ensure fairness. We evaluate four RAG variants: regular RAG Lewis et al. (2021), GraphRAG with local and global search Edge et al. (2024), and our proposed CausalRAG. Local search retrieves from raw document graphs, making it ideal for passage-specific queries, while global search relies on graph community summaries for broader context understanding. We use GPT-4o-mini as the base LLM for all models.
Datasets. We evaluate on publicly available research papers from various domains, ranging from 100 to around 17,000 tokens, reflecting typical user-uploaded documents. An LLM will generate $n$ grounded questions per document, ensuring they are explicitly answerable. We use the Ragas framework Es et al. (2023) and assess models on three key metrics: answer faithfulness, context precision, and context recall. Faithfulness measures factual consistency on a scale from 0 to 1, with higher scores indicating better alignment with reference documents.
5.2 Performance Comparison
To comprehensively evaluate the effectiveness of CausalRAG, we conducted experiments on a diverse dataset comprising abstracts from 100 randomly selected research papers using OpenAlexPriem et al. (2022). All evaluations were performed using the OpenAlex public academic dataset, which spans multiple disciplines, including applied mathematics, art history, library science, and psychology, ensuring a comprehensive assessment across different domains. For each document, we constructed graph and generated $n=5$ questions per paper for assessing retrieval and response quality. As mentioned, we then compared CausalRAG’s performance against three baselines—regular RAG, GraphRAG-Local, and GraphRAG-Global—across three key metrics: answer faithfulness, context recall, and context precision.
Figure 3 presents the comparative results. Overall, CausalRAG consistently outperforms all other models across the three evaluation metrics, demonstrating its ability to generate more factually accurate responses while maintaining both high recall and precision.
Answer Faithfulness. This metric measures the extent to which the generated response aligns with the reference information, ensuring factual correctness. As shown in Figure 3, regular RAG achieves relatively strong performance in this metric, suggesting that despite its reliance on purely semantic retrieval, it still retrieves some relevant context. However, GraphRAG-Local still slightly outperforms regular RAG by leveraging structured graph-based retrieval locally, which supports its grounded answer. Naturally, GraphRAG-Global CausalRAG supported by high-level community summary performs worse in this metric. Lastly, CausalRAG holds its answer faithfulness at a good level by retrieving causally grounded context, reducing hallucinations, and ensuring responses are not only relevant but also justified by retrieved evidence.
Context Recall. Context recall evaluates how much of the correct reference information is retrieved during the RAG process. As expected, regular RAG exhibits high recall, given its broad retrieval approach, which tends to maximize the inclusion of potentially relevant content. GraphRAG-Local and GraphRAG-Global show a slight reduction in recall due to their community-based summarization, which, while improving retrieving process, sacrifices some context diversity. CausalRAG, by contrast, strikes a balance between precision and recall and its recall is still higher than regular RAG. It is ensured by that retrieved information is causally relevant, avoiding excessive retrieval of loosely related but non-causal information.
Context Precision. Precision is where CausalRAG demonstrates the most significant advantage. While regular RAG achieves relatively low precision—often retrieving semantically similar yet contextually irrelevant content—GraphRAG-Local and GraphRAG-Global significantly improve upon this by leveraging graph structures to organize knowledge more effectively. However, both still struggle with retrieving causally relevant content and are lower than CausalRAG on this metric. We can see CausalRAG still maintain a great precision, as it inherently filters retrieved information through causal graph.
Figure 3: Performance comparison of Regular RAG, GraphRAG-Global, GraphRAG-Local, and CausalRAG on the OpenAlex dataset across three key metrics: answer faithfulness, context recall, and context precision.
5.3 Case Study
Figure 4: Case Study comparing retrieving process between different RAG systems.
Figure 5: Performance comparison of four RAGs across documents of varying lengths (Abstract, Introduction, Full Paper) based on mean values of answer faithfulness, context recall, and context precision.
We also conducted a case study to evaluate the four RAG variants on a research paper with long text, but testing them on its abstract, introduction, and full text respectively. This approach ensures content consistency while varying document length (ranging from 255 to 16,475 tokens), allowing us to assess the scalability of CausalRAG.
For each of the three text materials, we constructed a separate graph, tested $n=20$ questions, and evaluated all four RAG variants across three key metrics. As mentioned earlier, we employed an LLM-based approach for graph construction, and the validity of these graphs—built from a business school marketing paper—was further verified by experts.
This case study also examines the retrieval differences among regular RAG, GraphRAG, and our proposed CausalRAG (as shown in Figure 4). In this example, a user uploads a long paper and asks: "How do different combinations of influence tactics impact the likelihood of winning a sales contract?" Expert validation confirms a clear answer—salespeople use influence tactics such as compliance-based and internalization-based strategies to attract buyers’ attention, which inherently increases their chances of securing a contract.
Analyzing the retrieval process, we observe that regular RAG retrieves semantically related content but fails to capture key information, leading to a vague and uninformative response. GraphRAG correctly retrieves relevant nodes like attention, but its community report process dilutes their importance by incorporating extraneous nodes, introducing bias despite factual grounding. In contrast, CausalRAG retrieves the initial relevant nodes and expands upon them by identifying causal paths essential for answering the query, ensuring a precise and causally grounded response.
While this example clarifies CausalRAG’s internal process, our full experimental results (Figure 5) demonstrate its scalability. CausalRAG consistently outperforms other RAG variants across different document lengths. Additionally, regular RAG performs well on shorter documents, even surpassing GraphRAG. However, as document length increases—such as with full papers—GraphRAG overtakes regular RAG, aligning with our analytical analysis.
CausalRAG exhibits robustness across varying context lengths by preserving access to discrete information while leveraging causal estimation for grounded retrieval. More importantly, an additional advantage—only partially reflected here—is that with a larger step size $s$, CausalRAG can trace long and complex causal relationships hidden in lengthy documents, generating well-founded responses—a capability that is challenging for existing RAG systems.
5.4 Parameter Study
Figure 6: Parameter study showing how different parameter choices (k and s) affect model performance.
We also tested the impact of different parameter combinations of $k$ and $s$ on CausalRAG (as shown in Figure 6). Using the average of our evaluation metrics, we observe a consistent trend: the performance of CausalRAG improves as $k$ and $s$ increase. Specifically, the performance rises from 0.534 at $k=s=1$ to 0.824 at $k=s=5$, aligning with intuitive expectations.
Notably, the improvement is more pronounced when increasing $k$ from 1 to 3, suggesting that retrieving additional context enhances reasoning quality. However, when $k\geq 4$, performance gains become less significant, indicating possible saturation due to information redundancy. Similarly, while increasing $s$ generally leads to better results, its effect diminishes at higher values of $k$, where retrieval is already extensive.
These results suggest an optimal trade-off between performance and computational efficiency. While the highest values ($k=5,s=5$) yield the best results, moderate settings such as $k=3,s=3$ still achieve competitive performance with lower retrieval costs. Future work could explore adaptive strategies to adjust these parameters dynamically based on query complexity.
5.5 Conclusion and Future work
We introduced CausalRAG, a framework that integrates causality into RAG to enhance performance. Our analytical and experimental analysis identified key limitations of regular RAG—loss of contextual integrity, reliance on semantic similarity over causal relevance, and low precision. By leveraging causal graphs, CausalRAG retrieves causally grounded context, outperforming baseline RAGs across diverse text lengths and domains. Additionally, it improves LLM explainability, enhances answer groundedness, and reduces hallucinations in generative responses. Future work includes benchmarking CausalRAG on domain-specific tasks.
Limitations
While CausalRAG enhances retrieval effectiveness by integrating causality, it has certain limitations. First, constructing causal graphs from unstructured text relies on LLM-based extraction, which may introduce extra costs, particularly in complex or ambiguous cases. Second, the computational cost of expanding and analyzing causal paths increases with larger documents, potentially impacting retrieval efficiency in extreme cases with a large number of tokens.

### v2

@@
 1 Introduction
-The rapid advancements in large language models (LLMs) have revolutionized the field of natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024).
-However, their reliance on pre-trained knowledge limits their ability to integrate and reason over dynamically updated external information, particularly in knowledge intensive domains such as academic research.
-Retrieval-Augmented Generation (RAG) has emerged as a promising framework to address this limitation Lewis et al.
-(2021), combining retrieval mechanisms with generative capabilities to enhance contextual understanding and response quality.
-Recent research has focused on improving RAG along two primary directions: 1) enhancing retrieval efficiency and integration mechanisms by designing more adaptive and dynamic retrieval frameworks Gan et al.
+Large language models (LLMs) have transformed natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024).
+However, their reliance on static, pre-trained knowledge limits their ability to incorporate and reason over dynamically updated external information, particularly in knowledge-intensive domains.
+Retrieval-Augmented Generation (RAG) addresses this limitation by combining external retrieval with generative modeling to improve contextual understanding and response quality Lewis et al.
+(2021).
+Recent efforts to improve RAG focus on two fronts: 1) enhancing retrieval efficiency through adaptive and modular frameworks Gan et al.
 (2024); Ravuru et al.
 (2024); Zhang et al.
-(2024a); 2) improving the representation of external knowledge to facilitate retrieval and reasoning, with graph-based RAGs being a dominant approach Edge et al.
+(2024a); and 2) better structuring external knowledge, with graph-based RAGs emerging as a dominant approach Edge et al.
 (2024); Guo et al.
@@
 Despite these advancements, existing RAG architectures still face critical limitations that impact retrieval quality and response accuracy, primarily due to three key issues: 1) disruption of contextual integrity caused by the text chunking design; 2) reliance on semantic similarity rather than causal relevance for retrieval; and 3) a lack of accuracy in selecting truly relevant documents.
-Through a combination of theoretical analysis and empirical evaluation, we rethink the limitations of current RAG systems by introducing a novel perspective based on context recall and precision metrics.
-Our findings reveal that both regular and graph-based RAGs struggle not only to retrieve truly grounded context but also to accurately discern the relationship between retrieved content and the user query.
-We identify this fundamental issue as one primary reason why LLMs in RAG frameworks often generate seemingly relevant yet shallow responses that lack essential details.
-To address these gaps, we introduce CausalRAG, a novel RAG framework that integrates causal graphs to enhance retrieval accuracy and reasoning performance.
-Unlike regular and graph-based RAGs, CausalRAG constructs a causal graph from uploaded documents, preserving contextual relationships while capturing cause-effect dependencies.
-By ensuring that retrieved documents are both relevant and causally grounded, CausalRAG enables the generation of more contextually rich and causally detailed responses.
-This approach not only improves retrieval effectiveness but also mitigates hallucinations and enhances answer faithfulness.
-We evaluate CausalRAG on datasets from diverse domains and across varying context lengths, comparing its performance with regular RAG and GraphRAG, a graph-based RAG framework from Microsoft Edge et al.
-(2024).
-Our experiments assess performance across three key metrics: answer faithfulness, context recall, and context precision.
+Through theoretical and empirical analysis, we rethink the limitations of current RAG systems by introducing a novel perspective grounded in context recall and precision metrics.
+Using this lens, we find that both regular and graph-based RAGs often fail to retrieve causally grounded content or accurately align it with the user query.
+We identify this fundamental issue as a primary reason why LLMs in RAG frameworks frequently produce seemingly relevant yet shallow responses that lack essential details.
+To address these gaps, we propose CausalRAG, a novel RAG framework that integrates causal graphs to guide retrieval.
+By identifying cause–effect relationships within external knowledge, CausalRAG preserves contextual coherence and improves reasoning fidelity.
+This leads to more accurate and causally grounded responses, while reducing hallucinations and enhancing answer faithfulness.
+We evaluate CausalRAG across datasets and varying context lengths, comparing it with regular and other competitive graph-based RAGs over multiple metrics.
 Results demonstrate that CausalRAG achieves superior performance across different contexts.
@@
 The contributions of this work are threefold:
-  • We systematically identify the inherent limitations of RAG’s retrieval process through analytical and experimental study.
-More importantly, we uncover one major reason why LLMs in RAG tend to generate superficial, generalized answers that lack the grounded details expected by users.
-  • We propose CausalRAG, a framework that enhances both retrieval and generation quality by incorporating causality into the RAG, effectively addressing these limitations.
-  • Our work further mitigates hallucination issues and significantly improves the interpretability of AI systems.
-We summarize key findings and insights in both retrieval and generation process, contributing to research in RAG.
+  • We identify the root limitations of current RAG systems through analytical and empirical studies, revealing why LLMs often produce superficial and under-grounded responses.
+  • We propose CausalRAG, which improves retrieval and generation by grounding responses in causal structure.
+  • Our work further mitigates hallucination issues and enhance interpretability, offering new insights into the design of RAG systems.
 2 Related Work
@@
 (2021).
-Existing research primarily advances RAG along two key dimensions: 1) improving retrieval efficiency and integration mechanisms; 2) enhancing the representation of external knowledge to facilitate reasoning and interpretability.
+Existing research advances RAG in two main directions: 1) optimizing retrieval flow and interaction; and 2) structuring and utilizing external knowledge—particularly through knowledge graphs—for multi-stage reasoning.
 Optimizing retrieval flow and interaction.
-The first stream focuses on improving the interaction flow within the RAG system to enhance output quality.
-Approaches have introduced pre-retrieval, retrieval, and post-retrieval refinements to mitigate redundancy and computational overhead Wang et al.
+The first stream focuses on refining retrieval dynamics to improve output quality.
+Several methods introduce pre-, mid-, and post-retrieval enhancements to reduce redundancy and computation Wang et al.
 (2024).
-Modular RAG architectures further advance this by enabling iterative retrieval-generation cycles, allowing dynamic interactions between retrieval and content creation.
-For example, CAiRE-COVID Su et al.
-(2020) demonstrated the effectiveness of iterative retrieval in multi-document summarization, while some work Feng et al.
-(2023) extended this approach to multi-hop question answering.
-Recent innovations include METRAG Gan et al.
-(2024), which integrates LLM supervision to generate utility-driven retrieval processes, and RAFT Zhang et al.
-(2024a), which trains models to disregard distractor documents while improving citation accuracy through chain-of-thought reasoning.
-Structuring external knowledge for efficiency.
-The second direction explores improved methods for structuring external knowledge to achieve better retrieval efficiency.
-For example, GraphRAG Edge et al.
-(2024), proposed by Microsoft, treats external knowledge as interconnected nodes, capturing causal and thematic relationships to enhance retrieval depth and reasoning.
-LightRAG introduced a dual-level retrieval mechanism for incremental knowledge updates Guo et al.
+Modular RAG architectures enable iterative retrieval–generation loops for more adaptive workflows.
+For instance, CAiRE-COVID Su et al.
+(2020) demonstrated multi-document summarization with iterative retrieval, and others extended this to multi-hop QA Feng et al.
 (2023).
-More recently, Lazy GraphRAG has been developed as a continuous version of GraphRAG, deferring computationally expensive operations until query time and leveraging lightweight indexing techniques for efficiency Potts (2024).
-Despite these advancements, ensuring the quality and relevance of retrieved documents remains a significant concern in current RAG systems, as it directly impacts the coherence of the final response Gupta et al.
+Recent work like METRAG Gan et al.
+(2024) leverages LLM supervision for utility-driven retrieval, while RAFT Zhang et al.
+(2024a) trains models to disregard distractor documents through chain-of-thought reasoning.
+Structuring and utilizing external knowledge for multi-stage reasoning.
+The second stream focuses on representing and leveraging external knowledge—often via knowledge graphs—to support multi-level reasoning.
+GraphRAG Edge et al.
+(2024) represents documents as nodes and edges and summarizes subgraph communities for retrieval.
+LightRAG Guo et al.
+(2024) introduces dual-level retrieval and supports graph dynamic updates, while Lazy GraphRAG Potts (2024) further improves efficiency by delaying heavy computations until query time.
+More recent multi-stage RAGs formalize hierarchical reasoning across multiple stages.
+PolyRAG Chen et al.
+(2025) uses a knowledge pyramid for coarse-to-fine retrieval.
+KG2RAG Zhu et al.
+(2025) guides RAG with KG-guided chunk expansion process to enhance factual grounding and control.
+GenTKGQA Gao et al.
+(2024) performs two-stage retrieval and generation over temporal subgraphs.
+HippoRAG2 Gutiérrez et al.
+(2025) integrates PageRank ranking with deeper passage reasoning for better answer quality.
+However, most of these approaches focus primarily on matching and ranking information within the knowledge graph, with limited attention to the causal intent behind the user’s query.
+A key challenge in RAG remains ensuring that retrieved information is not only relevant but also coherently aligned with the user’s underlying reasoning needs and the generation objective Gupta et al.
 (2024).
 2.2 Causal Graphs and RAG
-The combination of causal graphs and RAG has emerged as a promising approach to enhancing knowledge retrieval and reasoning.
-As causality provides a structural understanding of dependencies within data, it enables more interpretable and reliable AI outputs Ma (2024).
-Existing research in this domain primarily advances causal discovery with RAG and LLMs.
-For instance, some work proposed an LLM-assisted breadth-first search (BFS) method for full causal graph discovery, significantly reducing time complexity Jiralerspong et al.
+Causal graph integration has emerged as a promising way to improve retrieval and reasoning.
+Causal structures support more interpretable and reliable outputs by modeling underlying dependencies Ma (2024).
+Prior work in this area primarily focuses on causal discovery using RAG and LLMs.
+For instance, LLM-assisted BFS has been used to accelerate full causal graph discovery Jiralerspong et al.
+(2024), while Corr2Cause evaluates LLMs’ ability to infer causation from correlation Jin et al.
 (2024).
-Additionally, some further introduced a correlation-to-causation inference (Corr2Cause) task to evaluate LLMs’ ability to infer causation from correlation, revealing their limitations in generalization across different datasets (Jin et al., 2024).
 Despite these advancements, most studies focus on utilizing RAG or LLMs for causal discovery or causal effect estimation Ma (2024); Kıcıman et al.
@@
 A few existing studies have touched upon this concept but differ in scope.
-One approach integrates causal graphs within the LLM architecture itself, structuring the transformer’s internal token processing using causality rather than enhancing RAG retrieval Chen et al.
-(2024b).
-Another employs causal graphs in RAG systems but focuses on the pre-retrieval stage and largely reduces the core process into a single embedding model without deeper exploration Samarajeewa et al.
+One integrates causality within the transformer architecture itself Chen et al.
+(2024), and another uses causal graphs only in pre-retrieval stages without deeper exploration Samarajeewa et al.
 (2024).
-GraphRAG is a well-regarded and influential work in this area, as it introduces a graph-based structure into the RAG system, leveraging graph community detection and summarization techniques for retrieval Edge et al.
-(2024).
-While it does not incorporate causality, it has significantly improved RAG performance.
-Therefore, we adopt GraphRAG as the baseline for our work.
-In the following sections, we first analyze the nature of regular RAG and grah-based RAGs from a novel perspective, identifying their inherent limitations.
-We then introduce a causal graph structure to address these gaps within the RAG system and present our framework—CausalRAG.
+GraphRAG Edge et al.
+(2024), while influential for its use of community detection and graph summarization, does not incorporate causality—nor do many recent multi-stage graph-based RAG frameworks.
+In the following sections, we first analyze the limitations of regular and graph-based RAGs using a new recall–precision lens. We then introduce a causal graph structure to address these gaps.
 Figure 1: Analytical and experimental studies reveal limitations in regular RAG and GraphRAG.
@@
 Context $\text{Context Recall}=\frac{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}{|R|}$ (1)
-Here $R$ is the reference set of all relevant reference documents.
-$C_{i}$ is the $i^{th}$ retrieved document.
+Here $R$ is the reference set of all relevant references.
+$C_{i}$ is the $i^{th}$ retrieved reference.
 $\mathbb{I}(C_{i}\in R)$ is an indicator function that returns 1 if $C_{i}$ belongs to the reference set $R$, otherwise 0.
@@
 Context $\text{Context Precision}=\frac{\sum_{i=1}^{N}\mathbb{I_{Q}}(C_{i}\in R)}{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}$ (2)
-Here $\mathbb{I_{Q}}(C_{i}\in R)$ is an indicator function that returns 1 if the context retrieved is causally related to user’s query, otherwise 0.
-Recall-Precision Perspective.
-Context recall measures how much of the correct contextual information can be retrieved from the user’s uploaded documents given a query.
-While RAG primarily focuses on retrieving related documents, adjusting retrieval parameters to increase the number of retrieved documents can generally lead to higher recall in reality.
-However, context precision presents a challenge.
-It measures the proportion of retrieved documents that are actually correct given the user query.
-As discussed earlier, RAG’s reliance on semantic similarity rather than causal relevance often leads to the retrieval of superficially similar but logically irrelevant content.
-In summary, while RAG can recall numerous answers from reference materials, the proportion of correct responses remains low, ultimately reducing its precision.
-This recall-precision perspective provides a new lens to see limitations of regular RAG.
-3.2 Rethinking Graph-based RAGs
-Applying this perspective, we can better understand why Graph-based RAGs serve as improved variants of RAG.
-By summarizing and ranking the importance of graph communities before retrieval, they largely enhance the quality of retrieved context, thereby improving context precision.
-However, it only partially addresses the identified limitations, as its summarization process does not entirely filter out irrelevant information.
-More critically, its reliance on community-based summarization for retrieval may adversely impact recall.
-Based on these analytical insights, we further conducted an experimental study to validate our analysis.
-Experimental study.
-As shown in Figure 1(b), we conducted an experimental study to empirically verify our analysis.
-Specifically, we employed an LLM to evaluate the performance of regular RAG and GraphRAG using the RAG evaluation framework Ragas Es et al.
-(2023).
-It takes as input the three key elements—user query, retrieved context, and reference—along with predefined metric definitions, and prompts the LLM to assign numerical ratings.
-This LLM-based evaluation has been widely used in recent RAG research and has also been adopted in RAG evaluations Samarajeewa et al.
-(2024); Edge et al.
-(2024).
-The results indicate that both global and local versions of GraphRAG achieve higher context precision compared to regular RAG.
-However, its recall performance remains unsatisfactory, even slightly lower than that of RAG, due to its graph community-based retrieval process.
-By combining our precision-recall-based analytical analysis with empirical findings, we clearly outline the inherent limitations of both RAG and its graph-based extensions.
-In the next section, we introduce our proposed framework, CausalRAG, designed to address these issues.
+Here $\mathbb{I_{Q}}(C_{i}\in R)$ is an indicator function that returns 1 if the context retrieved is causally related to the user’s query, otherwise 0.
 Figure 2: Overview of CausalRAG’s architecture.
@@
 A causal summary is generated and combined with the query to ensure grounded responses.
+Recall–precision perspective.
+Context recall measures the extent to which relevant contextual information can be retrieved from external knowledge given a query.
+In practice, increasing the number of retrieved documents typically improves recall in RAG systems.
+However, this semantic search–based approach often sacrifices precision—the proportion of retrieved content that is truly correct and directly relevant to the user query.
+For example, when asking, “How does this article define AI?”, regular RAG tends to retrieve all cited definitions that are semantically similar, even though only the author’s own definition is actually pertinent.
+This illustrates a core limitation of relying on semantic similarity rather than causal relevance: the retrieval of content that appears related but is logically irrelevant.
+More critically, low retrieval precision introduces systematic bias, diluting the accuracy and reliability of the retrieved context.
+In summary, while RAG can recall numerous answers from reference materials, the proportion of correct context remains low, ultimately reducing its precision.
+This recall-precision perspective provides a new lens to see the limitations of RAG frameworks.
+3.2 Rethinking Graph-based RAGs
+Applying this perspective, we can better understand why Graph-based RAGs serve as improved variants of RAG.
+By summarizing and ranking the importance of graph information before retrieval, they largely enhance the quality of retrieved context, thereby improving context precision.
+However, it only partially addresses the identified limitations, as its summarization process does not entirely filter out irrelevant information.
+More importantly, its reliance on subgraph summarization for retrieval may adversely impact recall by omitting less graph-central but still causally relevant context.
+To further examine these trade-offs, we conducted an experimental study to empirically validate these analytical insights.
+Experimental study.
+Figure 1(b) reports experimental results for Regular RAG and both variants of GraphRAG.
+The graph-based approaches markedly improve context precision, reflecting the additional reranking and subgraph-summary steps in their retrieval pipelines.
+This gain comes with a modest reduction in context recall, indicating that some relevant passages are pruned during summarization.
+These scores were obtained with the Ragas evaluation framework Es et al.
+(2023); more comprehensive experiments and implementation details are shown in Section 5.
+By combining our recall–precision-based analytical insights with empirical findings, we highlight the inherent limitations of both standard RAG and its graph-based extensions.
+Specifically, we show that relying on semantic similarity and subgraph summarization—rather than causal relationships—introduces trade-offs that often result in superficial and less accurate generation.
+We also present a case study in Section 5.3, which explicitly compares the retrieval processes across different RAG frameworks.
+In the following section, we introduce our proposed framework, CausalRAG, developed to address these limitations through causally grounded retrieval.
 4 Methodology
-In this section, we introduce our novel framework—CausalRAG—which integrates RAG with causality to overcome the limitations of existing RAG systems.
-Overall, CausalRAG utilizes a graph-based approach to represent uploaded documents, enhancing robustness against discontinuities caused by text chunking (as shown in Figure 2).
-More importantly, by expanding and tracing nodes in the graph through causal relationships, CausalRAG is able to retrieve a more diverse set of contextual information while maintaining causal groundedness.
-This enables CausalRAG to achieve both high recall and strong precision.
-We now discuss each step in detail.
+In this section, we introduce our proposed framework—CausalRAG—which integrates RAG with causality to overcome the limitations of existing RAG systems.
+Overall, CausalRAG constructs a text-based graph from uploaded documents and discovers causal paths among nodes to guide retrieval (Figure 2).
+By embedding the user query, matching relevant nodes, and expanding them along causally linked paths, the system generates a causal summary that serves as the retrieval context.
+This approach enables the framework to preserve contextual coherence and retrieve more targeted and causally relevant evidence.
+We now describe each step of the framework in detail.
 4.1 Indexing
-At the outset, once the system receives the user’s uploaded documents and query, we first index these inputs into our vector database.
-For the uploaded documents, we employ a text-based graph construction method to transform the text into a structured graph before storing the nodes and edges in the vector database.
-Specifically, we leverage an LLM to construct the graph following the approach by LangChain Chase (2022), where the LLM scans the text to identify graph nodes and determine relationships between them.
-Although LLM-based graph construction has shown strong performance Chen et al.
-(2024a) and is widely adopted in graph-based RAG research , we further validate our constructed graph using expert knowledge, which we elaborate on in the case study.
-After constructing and indexing the base graph, we embed the user query into the vector database, preparing it for subsequent search and matching.
-Notably, this indexing process occurs independently of query time, allowing for efficient retrieval during inference.
+At the outset, upon receiving the user’s uploaded documents and query, the system first indexes both inputs into a vector database.
+For the documents, we employ a text-based graph construction method that transforms unstructured text into a structured graph comprising nodes and edges.
+Specifically, we follow the approach proposed in LangChain Chase (2022), where an LLM parses the text to identify key entities or concepts as nodes and infers relationships between them as edges.
+Though this is a widely adopted approach in RAG research, we further validate the resulting graphs using expert knowledge, as discussed in the case study (Section 5.3).
+Once the graph is constructed, it is embedded and stored in the vector database, enabling efficient similarity-based retrieval.
+The user query is also embedded at this stage, preparing the system for subsequent matching.
+Importantly, this indexing process is performed offline and independently of query time, ensuring fast and scalable inference.
 4.2 Discovering and Estimating Causal Paths
-This is the first step during query time.
-We begin by matching the user query to nodes in the graph based on their embedding distance.
-The $k$ nodes with the smallest distances are selected, representing the most relevant information directly related to the query.
-Notably, $k$ is a tunable parameter—higher values retrieve more relevant information at the cost of increased computational complexity.
-After selecting the initial $k$ nodes, we expand the search along the base graph’s edges by a step size of $s$, thereby broadening the retrieved context.
-This step is crucial as it preserves causal and relational connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall.
-The parameter $s$ controls the depth of expansion, where higher values lead to more diverse information retrieval.
-Once the relevant nodes and edges are collected, we employ an LLM to identify and estimate causal paths within them, constructing a refined causal graph.
-As discussed, LLMs have demonstrated efficiency in discerning and analyzing causal relationships Zhang et al.
+At query time, we first match the user query to graph nodes using embedding distance and select the top-$k$ closest nodes.
+The parameter $k$ controls the retrieval breadth: higher values return more context but increase computational cost.
+We then expand these nodes along graph edges by a step size $s$, which allows the framework to preserve causally and relationally connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall.
+The parameter $s$ determines the expansion depth and diversity of retrieved information.
+Once the relevant nodes and edges are collected, we employ an LLM to identify and estimate causal paths within them, constructing a refined causal graph and generating the causal summary report (see LLM prompts in Appendix A).
+LLMs have demonstrated superiority in discerning and analyzing causal relationships Zhang et al.
 (2024b); Zhou et al.
+(2024) and this step ensures that CausalRAG prioritizes causally relevant information, improving precision.
+The resulting causal graph serves two purposes: (1) capturing long-range cause–effect dependencies often missed by semantic retrieval, particularly in longer texts; and (2) filtering out superficially related but causally irrelevant content, reducing hallucinations and improving faithfulness.
+Life Sciences Computing & Math Social Sciences Physics Other Total Tokens
+13.27% 14.29% 21.43% 14.29% 36.73% 21,285
+Table 1: Statistics of the dataset domain distribution and token lengths.
+Figure 3: Performance comparison of CausalRAG, regular RAG, and other graph-based RAGs across three key metrics: answer faithfulness, context recall, and context precision.
+4.3 Retrieving Context Causally
+After tracing and identifying key causal paths in the causal summary, we feed the result into the next stage and combine it with the user’s query (see LLM prompts in Appendix A).
+The causal summary prioritizes nodes that directly support the query, ensuring coherence and factual consistency while filtering out spurious information.
+This step reduces the risk of retrieving misleading yet superficially similar contexts.
+Finally, the combined prompt—including the causal summary—is used for generation, enabling CausalRAG to reason over major causal relationships rather than merely aggregating loosely related text spans.
+5 Experiment
+To evaluate the effectiveness of CausalRAG, we conduct a series of experiments comparing it with regular RAG and other competitive graph-based RAGs across multiple performance metrics to ensure a comprehensive assessment.
+In addition, we present a case study that explicitly compares the retrieval processes of different RAG variants, and a parameter study to further examine the behavior and performance of CausalRAG.
+5.1 Experimental Setup
+Baselines.
+We evaluate five RAG variants: Regular RAG Lewis et al.
+(2021), GraphRAG with both local and global search Edge et al.
+(2024), HippoRAG2 Gutiérrez et al.
+(2025), and our proposed CausalRAG.
+Regular RAG serves as a standard baseline, relying solely on semantic similarity for retrieval.
+GraphRAG is a widely recognized framework that leverages graph community summaries, and we include both of its modes: the local version retrieves from raw document graphs and is well-suited for passage-level queries, while the global version summarizes graph communities to support broader context understanding.
+HippoRAG2 is a recent and competitive method that enhances retrieval by selecting seed nodes and ranking filtered triples for generation.
+Datasets.
+While many open-domain QA benchmark datasets exist, most are designed for explicit fact retrieval (e.g., “When was Google founded?” “1998”) and target classic NLP tasks.
+These datasets often fall short in evaluating discourse-level understanding, such as querying the underlying ideas, logic, or narrative within a document—tasks that better reflect real-world needs.
+Although some reading comprehension datasets exist Rajpurkar et al.
+(2018), their answers are typically short entities and designed for NLP models rather than RAG systems.
+To more effectively evaluate RAGs in knowledge-intensive tasks, recent research calls for datasets that require higher-level discourse understanding, such as those based on podcasts, news articles, or Wikipedia Edge et al.
+(2024); Gutiérrez et al.
 (2025).
-This step ensures that CausalRAG prioritizes causally relevant information, improving precision.
-Furthermore, the derived causal graph serves two key purposes: 1) It preserves causally relevant information that traditional retrieval methods struggle to capture.
-More importantly, by adjusting the parameter $s$, this approach can capture long-range causal relationships within the text, particularly when the text is lengthy; 2) It filters out semantically related but causally irrelevant information.
-Without this filtering, responses may contain unnecessary or even hallucinated content, compromising answer faithfulness.
-4.3 Retrieving Context Causally
-After constructing the causal graph, we summarize the retrieved information and generate a causal summary.
-Notably, the input at this stage consists of information that is not only highly relevant but also causally grounded in the user’s query, ensuring greater validity.
-This approach contrasts with traditional retrieval methods, which often rely purely on semantic similarity and may retrieve contextually related yet causally irrelevant information.
-The causal summary is derived by tracing key causal paths within the graph, prioritizing nodes and relationships that contribute directly to answering the query.
-This ensures that the retrieved information maintains logical coherence and factual consistency while filtering out spurious or weakly related context.
-Additionally, by leveraging causal dependencies, our method reduces the risk of retrieving semantically similar but misleading evidence.
-Once the causal summary is generated, it is combined with the user query to construct a refined prompt for CausalRAG.
-This structured final input allows RAG to focus on reasoning through causal relationships rather than merely aggregating loosely related text spans.
-5 Experiment
-To evaluate the effectiveness of CausalRAG, we conduct a series of experiments comparing it with regular RAG and GraphRAG baselines.
-Our evaluation spans multiple datasets, retrieval settings, and performance metrics, ensuring a comprehensive analysis of retrieval quality and answer faithfulness.
-We systematically explore different parameter settings and retrieval strategies to assess their impact on model performance.
-5.1 Experimental Setup
-Parameters & Baselines.
-We set the default parameters for CausalRAG as $k=s=3$ and use the same $k$ value for GraphRAG’s community-based retrieval and regular RAG’s document retrieval to ensure fairness.
-We evaluate four RAG variants: regular RAG Lewis et al.
-(2021), GraphRAG with local and global search Edge et al.
-(2024), and our proposed CausalRAG.
-Local search retrieves from raw document graphs, making it ideal for passage-specific queries, while global search relies on graph community summaries for broader context understanding.
-We use GPT-4o-mini as the base LLM for all models.
-Datasets.
-We evaluate on publicly available research papers from various domains, ranging from 100 to around 17,000 tokens, reflecting typical user-uploaded documents.
-An LLM will generate $n$ grounded questions per document, ensuring they are explicitly answerable.
-We use the Ragas framework Es et al.
-(2023) and assess models on three key metrics: answer faithfulness, context precision, and context recall.
-Faithfulness measures factual consistency on a scale from 0 to 1, with higher scores indicating better alignment with reference documents.
+Following this direction, we use academic papers sampled from the OpenAlex dataset Priem et al.
+(2022a).
+Dataset statistics are provided in Table 1.
+For each document, we use an LLM to generate $n=5$ grounded questions, ensuring they are explicitly answerable (see Appendix B for examples).
+Metrics & implementation details.
+We use the Ragas evaluation framework Es et al.
+(2023) to assess all models on three metrics: answer faithfulness, context precision, and context recall.
+The definitions of context precision and context recall are provided in the previous section.
+Answer faithfulness measures factual consistency on a scale from 0 to 100, with higher scores indicating closer alignment with reference documents.
+We use GPT-4o-mini as the base LLM for all frameworks.
+We set the parameters $k=s=3$ for CausalRAG, and use the same $k$ value for GraphRAG’s community-based retrieval, Regular RAG’s document retrieval, and HippoRAG2’s triple-based retrieval to ensure a fair comparison.
+Figure 4: Case Study – A user uploads a long paper and asks a related question.
+This figure compares Regular RAG, GraphRAG, and CausalRAG by analyzing their retrieval processes.
+It highlights the drawbacks of semantic and graph-based retrieval and shows how causal reasoning in CausalRAG leads to more robust and precise results.
+Figure 5: Case Study – A follow-up experiment evaluates the RAGs in the previous case using three versions of the same paper.
+Graph-based methods improve with length, while CausalRAG remains consistently robust.
 5.2 Performance Comparison
-To comprehensively evaluate the effectiveness of CausalRAG, we conducted experiments on a diverse dataset comprising abstracts from 100 randomly selected research papers using OpenAlexPriem et al.
+Figure 3 presents the main experimental results comparing five RAG frameworks across three evaluation metrics: answer faithfulness, context recall, and context precision.
+The results reveal distinct patterns in how different retrieval strategies affect generation quality.
+Below, we summarize the key findings and underlying trade-offs.
+Causality delivers the most balanced and accurate retrieval.
+The model incorporating causal reasoning achieves the highest performance across all three metrics, with a context–precision score of $92.86$ and a faithfulness score of $78.00$, while maintaining competitive recall ($49.46$).
+By grounding retrieval in explicit causal paths, it filters out semantically similar but logically irrelevant content, thereby reducing superficial answers and improving the alignment and coherence between the retrieved context and the user query.
+Graph-level structuring improves precision but introduces a recall–coverage trade-off.
+GraphRAG’s two operating modes illustrate this tension.
+Local search, which queries the raw knowledge graph, yields strong precision ($89.18$) but discards many supporting passages, dropping recall to $41.54$.
+Global search restores some of that missing evidence by summarizing community subgraphs, yet the broader scope simultaneously re-introduces off-topic material, pulling precision down to $66.67$ and faithfulness below $55.3$.
+These results corroborate Edge et al.
+(2024)’s observation that finer-grained community answers excel at factual grounding, whereas higher-level summaries aid breadth but dilute specificity.
+Entity-centric multi-stage KGs narrow the gap but still overlook discourse context.
+HippoRAG2 augments Personalized PageRank with LLM-filtered triples and passage nodes, boosting precision to $73.72$ and faithfulness to $67.36$ without sacrificing recall relative to Regular RAG.
+Nevertheless, its reliance on entity extraction leaves many context-rich sentences unlinked to the query, a limitation they noted as concept–context trade-off and need for deeper contextualizationGutiérrez et al.
 (2025).
-All evaluations were performed using the OpenAlex public academic dataset, which spans multiple disciplines, including applied mathematics, art history, library science, and psychology, ensuring a comprehensive assessment across different domains.
-For each document, we constructed graph and generated $n=5$ questions per paper for assessing retrieval and response quality.
-As mentioned, we then compared CausalRAG’s performance against three baselines—regular RAG, GraphRAG-Local, and GraphRAG-Global—across three key metrics: answer faithfulness, context recall, and context precision.
-Figure 3 presents the comparative results.
-Overall, CausalRAG consistently outperforms all other models across the three evaluation metrics, demonstrating its ability to generate more factually accurate responses while maintaining both high recall and precision.
-Answer Faithfulness.
-This metric measures the extent to which the generated response aligns with the reference information, ensuring factual correctness.
-As shown in Figure 3, regular RAG achieves relatively strong performance in this metric, suggesting that despite its reliance on purely semantic retrieval, it still retrieves some relevant context.
-However, GraphRAG-Local still slightly outperforms regular RAG by leveraging structured graph-based retrieval locally, which supports its grounded answer.
-Naturally, GraphRAG-Global CausalRAG supported by high-level community summary performs worse in this metric.
-Lastly, CausalRAG holds its answer faithfulness at a good level by retrieving causally grounded context, reducing hallucinations, and ensuring responses are not only relevant but also justified by retrieved evidence.
-Context Recall.
-Context recall evaluates how much of the correct reference information is retrieved during the RAG process.
-As expected, regular RAG exhibits high recall, given its broad retrieval approach, which tends to maximize the inclusion of potentially relevant content.
-GraphRAG-Local and GraphRAG-Global show a slight reduction in recall due to their community-based summarization, which, while improving retrieving process, sacrifices some context diversity.
-CausalRAG, by contrast, strikes a balance between precision and recall and its recall is still higher than regular RAG.
-It is ensured by that retrieved information is causally relevant, avoiding excessive retrieval of loosely related but non-causal information.
-Context Precision.
-Precision is where CausalRAG demonstrates the most significant advantage.
-While regular RAG achieves relatively low precision—often retrieving semantically similar yet contextually irrelevant content—GraphRAG-Local and GraphRAG-Global significantly improve upon this by leveraging graph structures to organize knowledge more effectively.
-However, both still struggle with retrieving causally relevant content and are lower than CausalRAG on this metric.
-We can see CausalRAG still maintain a great precision, as it inherently filters retrieved information through causal graph.
-Figure 3: Performance comparison of Regular RAG, GraphRAG-Global, GraphRAG-Local, and CausalRAG on the OpenAlex dataset across three key metrics: answer faithfulness, context recall, and context precision.
+By contrast, CausalRAG’s path-based expansion integrates both entities and their causal explanations, yielding a retrieval set that is simultaneously broader in coverage and more tightly aligned with the question, thereby establishing a new possibility for knowledge-intensive RAG tasks.
 5.3 Case Study
-Figure 4: Case Study comparing retrieving process between different RAG systems.
-Figure 5: Performance comparison of four RAGs across documents of varying lengths (Abstract, Introduction, Full Paper) based on mean values of answer faithfulness, context recall, and context precision.
-We also conducted a case study to evaluate the four RAG variants on a research paper with long text, but testing them on its abstract, introduction, and full text respectively.
-This approach ensures content consistency while varying document length (ranging from 255 to 16,475 tokens), allowing us to assess the scalability of CausalRAG.
-For each of the three text materials, we constructed a separate graph, tested $n=20$ questions, and evaluated all four RAG variants across three key metrics.
-As mentioned earlier, we employed an LLM-based approach for graph construction, and the validity of these graphs—built from a business school marketing paper—was further verified by experts.
-This case study also examines the retrieval differences among regular RAG, GraphRAG, and our proposed CausalRAG (as shown in Figure 4).
-In this example, a user uploads a long paper and asks: "How do different combinations of influence tactics impact the likelihood of winning a sales contract?" Expert validation confirms a clear answer—salespeople use influence tactics such as compliance-based and internalization-based strategies to attract buyers’ attention, which inherently increases their chances of securing a contract.
-Analyzing the retrieval process, we observe that regular RAG retrieves semantically related content but fails to capture key information, leading to a vague and uninformative response.
-GraphRAG correctly retrieves relevant nodes like attention, but its community report process dilutes their importance by incorporating extraneous nodes, introducing bias despite factual grounding.
-In contrast, CausalRAG retrieves the initial relevant nodes and expands upon them by identifying causal paths essential for answering the query, ensuring a precise and causally grounded response.
-While this example clarifies CausalRAG’s internal process, our full experimental results (Figure 5) demonstrate its scalability.
-CausalRAG consistently outperforms other RAG variants across different document lengths.
-Additionally, regular RAG performs well on shorter documents, even surpassing GraphRAG.
-However, as document length increases—such as with full papers—GraphRAG overtakes regular RAG, aligning with our analytical analysis.
-CausalRAG exhibits robustness across varying context lengths by preserving access to discrete information while leveraging causal estimation for grounded retrieval.
-More importantly, an additional advantage—only partially reflected here—is that with a larger step size $s$, CausalRAG can trace long and complex causal relationships hidden in lengthy documents, generating well-founded responses—a capability that is challenging for existing RAG systems.
+Qualitative exemplar.
+Figure 4 traces the end-to-end behaviour of the three systems on a single question drawn from a marketing paper: “How do different combinations of influence tactics impact the likelihood of winning a sales contract?” The semantic baseline retrieves passages that mention “influence tactics” but stop short of establishing how those tactics translate into contract wins; its answer therefore echoes the query without providing the missing mechanism.
+Graph-based retrieval narrows the search space by clustering entities such as buyer attention and sales contract award, yet the community summary also injects peripheral phrases, leading the model to an over-generalised explanation that only partially matches the ground truth.
+By contrast, the causality-driven pipeline first aligns the query with a causal pathway—Influence Tactics $\rightarrow$ Buyer Attention $\rightarrow$ Contract Award.
+Because that pathway is preserved through retrieval and summarisation, the generated answer pinpoints the mediating role of buyer attention and specifies which tactic pairs are most effective, fully satisfying expert judgment.
+Length-controlled follow-up.
+To test whether these qualitative differences persist when context grows, we replicate the experiment on three versions of the same paper—abstract (250 tokens), introduction (1k), and full text (16k)—and score each system on the composite of faithfulness, precision, and recall (Figure 5).
+Graph-level methods benefit from additional material: their aggregate score climbs from 49.98/62.29 on the abstract to 76.37/84.27 on the full paper, confirming that community-based indexing scales gracefully with document length.
+Yet the causality-oriented approach remains consistently ahead, posting 72.43 on the abstract, 74.86 on the introduction, and 91.69 on the full text.
+This steadiness indicates that causal expansion recovers the right evidence even when local term overlap is sparse (short documents) and still filters noise when semantic matches proliferate (long documents).
+Emerging pattern.
+Taken together, the case study and its quantitative extension suggest a general rule: graph structure raises precision by adding relational cues, but explicit causal alignment is required to preserve both precision and recall across scales.
+Semantic search alone misses latent mechanisms; graph aggregation alone retains residual noise.
+Embedding causal constraints into retrieval therefore offers a principled path to robust performance on knowledge-intensive tasks, regardless of document length or query granularity.
 5.4 Parameter Study
@@
 5.5 Conclusion and Future Work
-We introduced CausalRAG, a framework that integrates causality into RAG to enhance performance.
-Our analytical and experimental analysis identified key limitations of regular RAG—loss of contextual integrity, reliance on semantic similarity over causal relevance, and low precision.
-By leveraging causal graphs, CausalRAG retrieves causally grounded context, outperforming baseline RAGs across diverse text lengths and domains.
-Additionally, it improves LLM explainability, enhances answer groundedness, and reduces hallucinations in generative responses.
-Future work includes benchmarking CausalRAG on domain-specific tasks.
+We introduced CausalRAG, a novel framework that integrates causal reasoning into retrieval-augmented generation to address core limitations of existing RAG systems.
+By leveraging causality, CausalRAG retrieves context that is not only relevant but also causally grounded—enhancing generation quality, reducing hallucinations, and improving answer fidelity.
+Future work may extend CausalRAG in several directions.
+A key area is testing scalability in long-context scenarios, which requires constructing and processing numerous high-dimensional graphs, each derived from large-scale documents spanning millions of tokens.
+While current practices often merge documents into a single graph or use smaller graphs due to data quality and computational constraints, this direction is critical for understanding RAG scalability.
+In addition to improving performance, evaluating efficiency and generation latency remains important for practical deployment.
 Limitations
-While CausalRAG enhances retrieval effectiveness by integrating causality, it has certain limitations.
-First, constructing causal graphs from unstructured text relies on LLM-based extraction, which may introduce extra costs, particularly in complex or ambiguous cases.
-Second, the computational cost of expanding and analyzing causal paths increases with larger documents, potentially impacting retrieval efficiency in extreme cases with a large number of tokens.
+While CausalRAG improves retrieval effectiveness through causal reasoning, it has certain limitations.
+First, the approach relies on the internal knowledge of LLMs to construct graphs and identify causal relationships.
+Although both capabilities have been actively studied, emerging domain-specific knowledge—such as in medicine or law—may still constrain its effectiveness in specialized contexts.
+Second, identifying causal paths during inference also requires additional LLM calls, introducing extra computational costs that may limit the performance in real-world deployments.
+Appendix
+Figure 7: LLM prompts for causal discovery and causal summary
+Appendix A LLM Prompts Used
+The full prompts used in the CausalRAG framework for causal discovery and causal summarization (Figure 7), as illustrated in Figure 2. The prompts are adapted from Edge et al. (2024).
+Appendix B OpenAlex Dataset Example
+This is an exmaple (see Table 2) in the dataset Priem et al. (2022b) along with generated questions.
+Dataset OpenAlex Priem et al. (2022b)
+Discipline Atomic Physics
+Author Thom H. Dunning (1989)
+Text In the past, basis sets for correlated molecular calculations were largely taken from single-configuration calculations.
+Recently, Almlöf, Taylor, and co-workers showed that atomic natural orbitals (ANOs) provide an excellent description of correlation effects.
+Here we report a careful oxygen-atom study establishing that compact primitive Gaussian functions effectively describe correlation when their exponents are optimized.
+Tests on oxygen-containing molecules indicate these functions perform as well as the ANO sets of Almlöf and Taylor.
+Guided by the oxygen results, basis sets were developed for all first-row atoms (B–Ne) and hydrogen.
+Incremental energy lowerings due to correlating functions fall into distinct groups, leading to the concept of consistent sets.
+The most accurate set, [5s 4p 3d 2f 1g], consistently yields 99% of the correlation energy obtained with the next larger set, even though the latter contains 50% more primitives and twice as many polarization functions.
+For boron, this set recovers 94–97% of the total (HF+1+2) correlation energy.
+Questions 1.
+What recent findings support the use of atomic natural orbitals in molecular calculations?
+2.
+How do compact primitive Gaussian functions contribute to describing correlation effects in oxygen?
+3.
+Why is exponent optimization important in Gaussian-based calculations?
+4.
+How do the new first-row basis sets compare in energy lowering from correlation effects?
+5.
+What accuracy (%) is achieved for boron with the most compact set, and how does this relate to the number of polarization functions?
+Table 2: Example document metadata, full-text excerpt, and evaluation questions used in our study.

### v3

@@
 We evaluate CausalRAG against regular RAG and graph-based RAG approaches, demonstrating its superiority across several metrics.
-Our findings suggest that grounding retrieval in causal reasoning provides a promising approach to knowledge-intensive tasks.
+Our findings suggest that grounding retrieval in causal reasoning provides a promising approach to knowledge-intensive tasks [^1] Code is publicly available at https://github.com/Pwnb/CausalRAG..
 CausalRAG: Integrating Causal Graphs into Retrieval-Augmented Generation
@@
 1 Introduction
-Large language models (LLMs) have transformed natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024).
-However, their reliance on static, pre-trained knowledge limits their ability to incorporate and reason over dynamically updated external information, particularly in knowledge-intensive domains.
-Retrieval-Augmented Generation (RAG) addresses this limitation by combining external retrieval with generative modeling to improve contextual understanding and response quality Lewis et al.
-(2021).
-Recent efforts to improve RAG focus on two fronts: 1) enhancing retrieval efficiency through adaptive and modular frameworks Gan et al.
+The rapid advancements in large language models (LLMs) have revolutionized the field of natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024).
+However, their reliance on pre-trained knowledge limits their ability to integrate and reason over dynamically updated external information, particularly in knowledge intensive domains such as academic research.
+Retrieval-Augmented Generation (RAG) has emerged as a promising framework to address this limitation Lewis et al.
+(2021), combining retrieval mechanisms with generative capabilities to enhance contextual understanding and response quality.
+Recent research has focused on improving RAG along two primary directions: 1) enhancing retrieval efficiency and integration mechanisms by designing more adaptive and dynamic retrieval frameworks Gan et al.
 (2024); Ravuru et al.
 (2024); Zhang et al.
-(2024a); and 2) better structuring external knowledge, with graph-based RAGs emerging as a dominant approach Edge et al.
+(2024a); 2) improving the representation of external knowledge to facilitate retrieval and reasoning, with graph-based RAGs being a dominant approach Edge et al.
 (2024); Guo et al.
@@
 Despite these advancements, existing RAG architectures still face critical limitations that impact retrieval quality and response accuracy, primarily due to three key issues: 1) disruption of contextual integrity caused by the text chunking design; 2) reliance on semantic similarity rather than causal relevance for retrieval; and 3) a lack of accuracy in selecting truly relevant documents.
-Through theoretical and empirical analysis, we rethink the limitations of current RAG systems by introducing a novel perspective grounded in context recall and precision metrics.
-Using this lens, we find that both regular and graph-based RAGs often fail to retrieve causally grounded content or accurately align it with the user query.
-We identify this fundamental issue as a primary reason why LLMs in RAG frameworks frequently produce seemingly relevant yet shallow responses that lack essential details.
-To address these gaps, we propose CausalRAG, a novel RAG framework that integrates causal graphs to guide retrieval.
-By identifying cause–effect relationships within external knowledge, CausalRAG preserves contextual coherence and improves reasoning fidelity.
-This leads to more accurate and causally grounded responses, while reducing hallucinations and enhancing answer faithfulness.
-We evaluate CausalRAG across datasets and varying context lengths, comparing it with regular and other competitive graph-based RAGs over multiple metrics.
+Through a combination of theoretical analysis and empirical evaluation, we rethink the limitations of current RAG systems by introducing a novel perspective based on context recall and precision metrics.
+Our findings reveal that both regular and graph-based RAGs struggle not only to retrieve truly grounded context but also to accurately discern the relationship between retrieved content and the user query.
+We identify this fundamental issue as one primary reason why LLMs in RAG frameworks often generate seemingly relevant yet shallow responses that lack essential details.
+To address these gaps, we introduce CausalRAG, a novel RAG framework that integrates causal graphs to enhance retrieval accuracy and reasoning performance.
+Unlike regular and graph-based RAGs, CausalRAG explicitly identifies causal relationships within external knowledge, preserving contextual coherence while capturing underlying cause–effect dependencies.
+By ensuring that retrieved documents are both relevant and causally grounded, CausalRAG enables the generation of more contextually rich and causally detailed responses.
+This approach not only improves retrieval effectiveness but also mitigates hallucinations and enhances answer faithfulness.
+We evaluate CausalRAG on datasets from diverse domains and across varying context lengths, comparing its performance with regular RAG and other graph-based RAG frameworks.
+Our experiments assess performance across three key metrics: answer faithfulness, context recall, and context precision.
 Results demonstrate that CausalRAG achieves superior performance across different contexts.
@@
 The contributions of this work are threefold:
-  • We identify the root limitations of current RAG systems through analytical and empirical studies, revealing why LLMs often produce superficial and under-grounded responses.
-  • We propose CausalRAG, which improves retrieval and generation by grounding responses in causal structure.
-  • Our work further mitigates hallucination issues and enhance interpretability, offering new insights into the design of RAG systems.
+  • We systematically identify the inherent limitations of RAG’s retrieval process through analytical and experimental study.
+We uncovered the major reason why LLMs in RAG tend to generate superficial, generalized answers that lack the grounded details expected by users.
+  • We propose CausalRAG, a framework that enhances both retrieval and generation quality by incorporating causality into the RAG, effectively addressing these limitations.
+  • Our work further mitigates hallucination issues and significantly improves the interpretability of AI systems.
+We summarize key findings and insights in both retrieval and generation process, contributing to RAG research.
 2 Related Work
@@
 (2021).
-Existing research advances RAG in two main directions: 1) optimizing retrieval flow and interaction; and 2) structuring and utilizing external knowledge—particularly through knowledge graphs—for multi-stage reasoning.
+Existing research primarily advances RAG along two key dimensions: 1) improving retrieval efficiency and integration mechanisms; 2) enhancing the representation and multi-stage utilization of external knowledge—particularly through knowledge graphs—to support deeper reasoning and generation.
 Optimizing Retrieval Flow and Interaction.
-The first stream focuses on refining retrieval dynamics to improve output quality.
-Several methods introduce pre-, mid-, and post-retrieval enhancements to reduce redundancy and computation Wang et al.
+The first stream focuses on improving the interaction flow within the RAG system to enhance output quality.
+Approaches have introduced pre-retrieval, retrieval, and post-retrieval refinements to mitigate redundancy and computational overhead Wang et al.
 (2024).
-Modular RAG architectures enable iterative retrieval–generation loops for more adaptive workflows.
-For instance, CAiRE-COVID Su et al.
-(2020) demonstrated multi-document summarization with iterative retrieval, and others extended this to multi-hop QA Feng et al.
-(2023).
-Recent work like METRAG Gan et al.
-(2024) leverages LLM supervision for utility-driven retrieval, while RAFT Zhang et al.
-(2024a) trains models to disregard distractor documents through chain-of-thought reasoning.
+Modular RAG architectures further advance this by enabling iterative retrieval-generation cycles, allowing dynamic interactions between retrieval and content creation.
+For example, CAiRE-COVID Su et al.
+(2020) demonstrated the effectiveness of iterative retrieval in multi-document summarization, while some work Feng et al.
+(2023) extended this approach to multi-hop question answering.
+Recent innovations include METRAG Gan et al.
+(2024), which integrates LLM supervision to generate utility-driven retrieval processes, and RAFT Zhang et al.
+(2024a), which trains models to disregard distractor documents while improving citation accuracy through chain-of-thought reasoning.
 Structuring and Utilizing External Knowledge for Multi-stage Reasoning.
-The second stream focuses on representing and leveraging external knowledge—often via knowledge graphs—to support multi-level reasoning.
+The second stream focuses on representing and processing external knowledge—often in the form of knowledge graphs—to improve retrieval depth and reasoning accuracy.
 GraphRAG Edge et al.
-(2024) represents documents as nodes and edges and summarizes subgraph communities for retrieval.
-LightRAG Guo et al.
-(2024) introduces dual-level retrieval and supports graph dynamic updates, while Lazy GraphRAG Potts (2024) further improves efficiency by delaying heavy computations until query time.
-More recent multi-stage RAGs formalize hierarchical reasoning across multiple stages.
+(2024) treats documents as interconnected nodes to capture thematic and causal links, while LightRAG Guo et al.
+(2024) adopts a lightweight, dual-level retrieval approach for dynamic updates.
+Lazy GraphRAG Potts (2024) further improves efficiency by delaying heavy computations until query time.
+More recently, several multi-stage RAG frameworks have emerged that explicitly structure the retrieval and generation process across multiple levels.
 PolyRAG Chen et al.
-(2025) uses a knowledge pyramid for coarse-to-fine retrieval.
+(2025) builds a hierarchical knowledge pyramid, allowing coarse-to-fine retrieval over different abstraction layers.
 KG2RAG Zhu et al.
@@
 GenTKGQA Gao et al.
-(2024) performs two-stage retrieval and generation over temporal subgraphs.
-HippoRAG2 Gutiérrez et al.
-(2025) integrates PageRank ranking with deeper passage reasoning for better answer quality.
+(2024) employs a two-stage pipeline to first retrieve temporal subgraphs and then generate based on virtual knowledge indicators.
+Among them, HippoRAG2 Gutiérrez et al.
+(2025) focuses on combining PageRank algorithm with deeper passage integration, improving both knowledge retention and answer quality.
 However, most of these approaches focus primarily on matching and ranking information within the knowledge graph, with limited attention to the causal intent behind the user’s query.
@@
 (2024).
+Figure 1: Analytical and experimental studies reveal limitations in regular RAG and GraphRAG.
+(a) identifies three key retrieval and generation issues in regular RAG; (b) evaluates RAG via context precision and recall, showing regular RAG excels in recall but lacks precision.
+GraphRAG improves precision but trades off some recall.
 2.2 Causal Graphs and RAG
-Causal graph integration has emerged as a promising way to improve retrieval and reasoning.
-Causal structures support more interpretable and reliable outputs by modeling underlying dependencies Ma (2024).
-Prior work in this area primarily focuses on causal discovery using RAG and LLMs.
-For instance, LLM-assisted BFS has been used to accelerate full causal graph discovery Jiralerspong et al.
-(2024), while Corr2Cause evaluates LLMs’ ability to infer causation from correlation Jin et al.
+The combination of causal graphs and RAG has emerged as a promising approach to enhancing knowledge retrieval and reasoning.
+As causality provides a structural understanding of dependencies within data, it enables more interpretable and reliable AI outputs Ma (2024).
+Existing research in this domain primarily advances causal discovery with RAG and LLMs.
+For instance, some work proposed an LLM-assisted breadth-first search (BFS) method for full causal graph discovery, significantly reducing time complexity Jiralerspong et al.
 (2024).
+Additionally, some further introduced a correlation-to-causation inference (Corr2Cause) task to evaluate LLMs’ ability to infer causation from correlation, revealing their limitations in generalization across different datasets (Jin et al., 2024).
 Despite these advancements, most studies focus on utilizing RAG or LLMs for causal discovery or causal effect estimation Ma (2024); Kıcıman et al.
@@
 A few existing studies have touched upon this concept but differ in scope.
-One integrates causality within the transformer architecture itself Chen et al.
-(2024), and another uses causal graphs only in pre-retrieval stages without deeper exploration Samarajeewa et al.
+One approach integrates causal graphs within the LLM architecture itself, structuring the transformer’s internal token processing using causality rather than enhancing RAG retrieval Chen et al.
+(2024).
+Another employs causal graphs in RAG systems but focuses on the pre-retrieval stage and largely reduces the core process into a single embedding model without deeper exploration Samarajeewa et al.
 (2024).
@@
 (2024), while influential for its use of community detection and graph summarization, does not incorporate causality—nor do many recent multi-stage graph-based RAG frameworks.
-In the following sections, we first analyze the limitations of regular and graph-based RAGs using a new recall–precision lens. We then introduce a causal graph structure to address these gaps.
-Figure 1: Analytical and experimental studies reveal limitations in regular RAG and GraphRAG.
-(a) identifies three key retrieval and generation issues in regular RAG; (b) evaluates RAG via context precision and recall, showing regular RAG excels in recall but lacks precision.
-GraphRAG improves precision but trades off some recall.
+In the following sections, we first analyze the nature of regular RAG and grah-based RAGs from a novel perspective, identifying their inherent limitations.
+We then introduce a causal graph structure to address these gaps.
 3 Why Regular RAG Fails in Providing Accurate Responses
@@
 To formalize this issue, we used two key metrics: context recall and context precision, defined as follows:
-Context $\text{Context Recall}=\frac{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}{|R|}$ (1)
+$\text{Context Recall}=\frac{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}{|R|}$ (1)
 Here $R$ is the reference set of all relevant references.
@@
 It should always return 1 if no hallucination occurs in the LLM.
-Context $\text{Context Precision}=\frac{\sum_{i=1}^{N}\mathbb{I_{Q}}(C_{i}\in R)}{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}$ (2)
+$\text{Context Precision}=\frac{\sum_{i=1}^{N}\mathbb{I_{Q}}(C_{i}\in R)}{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}$ (2)
 Here $\mathbb{I_{Q}}(C_{i}\in R)$ is an indicator function that returns 1 if the context retrieved is causally related to the user’s query, otherwise 0.
-Figure 2: Overview of CausalRAG’s architecture.
-Documents are indexed as graphs, and queries retrieve causally related nodes.
-A causal summary is generated and combined with the query to ensure grounded responses.
 Recall–precision Perspective.
@@
 Figure 1(b) reports experimental results for Regular RAG and both variants of GraphRAG.
-The graph-based approaches markedly improve context precision, reflecting the additional reranking and subgraph-summary steps in their retrieval pipelines.
+The graph‑based approaches markedly improve context precision, reflecting the additional reranking and subgraph‑summary steps in their retrieval pipelines.
 This gain comes with a modest reduction in context recall, indicating that some relevant passages are pruned during summarization.
@@
 In the following section, we introduce our proposed framework, CausalRAG, developed to address these limitations through causally grounded retrieval.
+Figure 2: Overview of CausalRAG’s architecture.
+Documents are indexed as graphs, and queries retrieve causally related nodes.
+A causal summary is generated and combined with the query to ensure grounded responses.
 4 Methodology
@@
 4.2 Discovering and Estimating Causal Paths
-At query time, we first match the user query to graph nodes using embedding distance and select the top-$k$ closest nodes.
-The parameter $k$ controls the retrieval breadth: higher values return more context but increase computational cost.
-We then expand these nodes along graph edges by a step size $s$, which allows the framework to preserve causally and relationally connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall.
-The parameter $s$ determines the expansion depth and diversity of retrieved information.
+This is the first step during query time.
+We begin by matching the user query to nodes in the graph based on their embedding distance.
+The $k$ nodes with the smallest distances are selected, representing the most relevant information directly related to the query.
+Notably, $k$ is a tunable parameter—higher values retrieve more relevant information at the cost of increased computational complexity.
+After selecting the initial $k$ nodes, we expand the search along the base graph’s edges by a step size of $s$, thereby broadening the retrieved context.
+This step is crucial as it preserves causal and relational connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall.
+The parameter $s$ controls the depth of expansion, where higher values lead to more diverse information retrieval.
 Once the relevant nodes and edges are collected, we employ an LLM to identify and estimate causal paths within them, constructing a refined causal graph and generating the causal summary report (see LLM prompts in Appendix A).
@@
 (2024) and this step ensures that CausalRAG prioritizes causally relevant information, improving precision.
-The resulting causal graph serves two purposes: (1) capturing long-range cause–effect dependencies often missed by semantic retrieval, particularly in longer texts; and (2) filtering out superficially related but causally irrelevant content, reducing hallucinations and improving faithfulness.
+Furthermore, the derived causal graph serves two key purposes: 1) It preserves causally relevant information that traditional retrieval methods struggle to capture.
+More importantly, by adjusting the parameter $s$, this approach can capture long-range causal relationships within the text, particularly when the text is lengthy; 2) It filters out semantically related but causally irrelevant information.
+Without this filtering, responses may contain unnecessary or even hallucinated content, compromising answer faithfulness.
 Life Sciences Computing & Math Social Sciences Physics Other Total Tokens
@@
 4.3 Retrieving Context Causally
-After tracing and identifying key causal paths in the causal summary, we feed the result into the next stage and combine it with the user’s query (see LLM prompts in Appendix A).
-The causal summary prioritizes nodes that directly support the query, ensuring coherence and factual consistency while filtering out spurious information.
-This step reduces the risk of retrieving misleading yet superficially similar contexts.
-Finally, the combined prompt—including the causal summary—is used for generation, enabling CausalRAG to reason over major causal relationships rather than merely aggregating loosely related text spans.
+After constructing the causal graph, we summarize the retrieved information and generate a causal summary (see LLM prompts in Appendix A).
+Notably, the input at this stage consists of information that is not only highly relevant but also causally grounded in the user’s query, ensuring greater validity.
+This approach contrasts with traditional retrieval methods, which often rely purely on semantic similarity and may retrieve contextually related yet causally irrelevant information.
+The causal summary is derived by tracing key causal paths within the graph, prioritizing nodes and relationships that contribute directly to answering the query.
+This ensures that the retrieved information maintains logical coherence and factual consistency while filtering out spurious or weakly related context.
+Additionally, by leveraging causal dependencies, our method reduces the risk of retrieving semantically similar but misleading evidence.
+Once the causal summary is generated, it is combined with the user query to construct a refined prompt for CausalRAG.
+This structured final input allows RAG to focus on reasoning through causal relationships rather than merely aggregating loosely related text spans.
 5 Experiment
@@
 Datasets.
-While many open-domain QA benchmark datasets exist, most are designed for explicit fact retrieval (e.g., “When was Google founded?” “1998”) and target classic NLP tasks.
+While many open-domain QA benchmark datasets exist, most are designed for explicit fact retrieval (e.g., “When was Google founded?” “1998.”) and target classic NLP tasks.
 These datasets often fall short in evaluating discourse-level understanding, such as querying the underlying ideas, logic, or narrative within a document—tasks that better reflect real-world needs.
@@
 For each document, we use an LLM to generate $n=5$ grounded questions, ensuring they are explicitly answerable (see Appendix B for examples).
-Metrics & implementation details.
+Metrics and Implementation Details.
 We use the Ragas evaluation framework Es et al.
@@
 Causality delivers the most balanced and accurate retrieval.
-The model incorporating causal reasoning achieves the highest performance across all three metrics, with a context–precision score of $92.86$ and a faithfulness score of $78.00$, while maintaining competitive recall ($49.46$).
-By grounding retrieval in explicit causal paths, it filters out semantically similar but logically irrelevant content, thereby reducing superficial answers and improving the alignment and coherence between the retrieved context and the user query.
-Graph-level structuring improves precision but introduces a recall–coverage trade-off.
-GraphRAG’s two operating modes illustrate this tension.
-Local search, which queries the raw knowledge graph, yields strong precision ($89.18$) but discards many supporting passages, dropping recall to $41.54$.
-Global search restores some of that missing evidence by summarizing community subgraphs, yet the broader scope simultaneously re-introduces off-topic material, pulling precision down to $66.67$ and faithfulness below $55.3$.
+The experimental results, as depicted in Figure 3, reveal distinct strengths and weaknesses among the evaluated RAG frameworks, ultimately highlighting the advantages of integrating causal reasoning.
+CausalRAG consistently achieves the most balanced and superior performance, leading in answer faithfulness ($78.00$), context precision ($92.86$), and maintaining competitive context recall ($49.46$).
+This suggests that by explicitly tracing and retrieving along causal paths, CausalRAG effectively mitigates the common RAG pitfalls of relying purely on semantic similarity and grappling with contextual integrity.
+Graph‑level structuring improves precision but introduces a recall–coverage trade‑off.
+Graph-based RAG approaches, while an improvement over regular RAG in certain aspects, demonstrate a clear trade-off.
+GraphRAG-Local excels in context precision ($89.18$) but suffers from a low context recall ($41.54$).
+This indicates that while its direct querying of the raw knowledge graph yields highly relevant local information, it often discards broader supporting passages.
+The global variant, GraphRAG-Global, attempts to address this by summarizing community subgraphs, thereby improving recall ($47.22$).
+However, this comes at a significant cost to precision ($66.67$) and faithfulness ($55.27$).
 These results corroborate Edge et al.
-(2024)’s observation that finer-grained community answers excel at factual grounding, whereas higher-level summaries aid breadth but dilute specificity.
-Entity-centric multi-stage KGs narrow the gap but still overlook discourse context.
-HippoRAG2 augments Personalized PageRank with LLM-filtered triples and passage nodes, boosting precision to $73.72$ and faithfulness to $67.36$ without sacrificing recall relative to Regular RAG.
-Nevertheless, its reliance on entity extraction leaves many context-rich sentences unlinked to the query, a limitation they noted as concept–context trade-off and need for deeper contextualizationGutiérrez et al.
+(2024)’s observation that finer‑grained community answers excel at factual grounding, whereas higher‑level summaries aid breadth but dilute specificity.
+In contrast, CausalRAG manages to maintain high precision while achieving better recall through its targeted causal path selection.
+Entity‑centric multi‑stage KGs narrow the gap but still overlook discourse context.
+HippoRAG2 presents a strong performance, particularly in answer faithfulness ($67.36$) and context precision ($73.72$) when compared to Regular RAG and GraphRAG-Global, with a recall of $47.22$.
+Its methodology, which augments Personalized PageRank with LLM-filtered triples and passage nodes, clearly enhances retrieval quality.
+Nevertheless, its reliance on entity extraction leaves many context‑rich sentences unlinked to the query, a limitation they noted as concept–context trade‑off and need for deeper contextualizationGutiérrez et al.
 (2025).
-By contrast, CausalRAG’s path-based expansion integrates both entities and their causal explanations, yielding a retrieval set that is simultaneously broader in coverage and more tightly aligned with the question, thereby establishing a new possibility for knowledge-intensive RAG tasks.
+Then we found CausalRAG surpasses it across all metrics.
+By comparison, CausalRAG’s causal path expansion allows for the integration of both entities and their underlying causal explanations.
+This approach yields a retrieval set that is not only broad in coverage but also more tightly aligned with the user’s query, leading to more accurate and faithful responses by ensuring the retrieved context is causally pertinent rather than just semantically adjacent.
 5.3 Case Study
 Qualitative Exemplar.
-Figure 4 traces the end-to-end behaviour of the three systems on a single question drawn from a marketing paper: “How do different combinations of influence tactics impact the likelihood of winning a sales contract?” The semantic baseline retrieves passages that mention “influence tactics” but stop short of establishing how those tactics translate into contract wins; its answer therefore echoes the query without providing the missing mechanism.
-Graph-based retrieval narrows the search space by clustering entities such as buyer attention and sales contract award, yet the community summary also injects peripheral phrases, leading the model to an over-generalised explanation that only partially matches the ground truth.
-By contrast, the causality-driven pipeline first aligns the query with a causal pathway—Influence Tactics $\rightarrow$ Buyer Attention $\rightarrow$ Contract Award.
+Figure 4 traces the end‑to‑end behaviour of the three systems on a single question drawn from a marketing paper Singh et al.
+(2020): “How do different combinations of influence tactics impact the likelihood of winning a sales contract?” The semantic baseline retrieves passages that mention “influence tactics” but stop short of establishing how those tactics translate into contract wins; its answer therefore echoes the query without providing the missing mechanism.
+Graph‑based retrieval narrows the search space by clustering entities such as buyer attention and sales contract award, yet the community summary also injects peripheral phrases, leading the model to an over‑generalised explanation that only partially matches the ground truth.
+By contrast, the causality‑driven pipeline first aligns the query with a causal pathway—Influence Tactics $\rightarrow$ Buyer Attention $\rightarrow$ Contract Award.
 Because that pathway is preserved through retrieval and summarisation, the generated answer pinpoints the mediating role of buyer attention and specifies which tactic pairs are most effective, fully satisfying expert judgment.
-Length-controlled follow-up.
+Length‑controlled Follow‑up.
 To test whether these qualitative differences persist when context grows, we replicate the experiment on three versions of the same paper—abstract (250 tokens), introduction (1k), and full text (16k)—and score each system on the composite of faithfulness, precision, and recall (Figure 5).
-Graph-level methods benefit from additional material: their aggregate score climbs from 49.98/62.29 on the abstract to 76.37/84.27 on the full paper, confirming that community-based indexing scales gracefully with document length.
-Yet the causality-oriented approach remains consistently ahead, posting 72.43 on the abstract, 74.86 on the introduction, and 91.69 on the full text.
+Graph‑level methods benefit from additional material: their aggregate score climbs from 49.98/62.29 on the abstract to 76.37/84.27 on the full paper, confirming that community‑based indexing scales gracefully with document length.
+Yet the causality‑oriented approach remains consistently ahead, posting 72.43 on the abstract, 74.86 on the introduction, and 91.69 on the full text.
 This steadiness indicates that causal expansion recovers the right evidence even when local term overlap is sparse (short documents) and still filters noise when semantic matches proliferate (long documents).
@@
 Semantic search alone misses latent mechanisms; graph aggregation alone retains residual noise.
-Embedding causal constraints into retrieval therefore offers a principled path to robust performance on knowledge-intensive tasks, regardless of document length or query granularity.
+Embedding causal constraints into retrieval therefore offers a principled path to robust performance on knowledge‑intensive tasks, regardless of document length or query granularity.
 5.4 Parameter Study
@@
 5.5 Conclusion and Future Work
-We introduced CausalRAG, a novel framework that integrates causal reasoning into retrieval-augmented generation to address core limitations of existing RAG systems.
-By leveraging causality, CausalRAG retrieves context that is not only relevant but also causally grounded—enhancing generation quality, reducing hallucinations, and improving answer fidelity.
-Future work may extend CausalRAG in several directions.
-A key area is testing scalability in long-context scenarios, which requires constructing and processing numerous high-dimensional graphs, each derived from large-scale documents spanning millions of tokens.
-While current practices often merge documents into a single graph or use smaller graphs due to data quality and computational constraints, this direction is critical for understanding RAG scalability.
-In addition to improving performance, evaluating efficiency and generation latency remains important for practical deployment.
+We introduced CausalRAG, a novel framework that integrates causal reasoning into retrieval-augmented generation to address key limitations of existing RAG systems.
+Through theoretical analysis and empirical validation, we demonstrated that regular RAGs suffer from disrupted contextual integrity, semantic-over-causal retrieval, and context accuracy trade-offs.
+By leveraging causal graphs, CausalRAG retrieves context that is not only relevant but causally grounded—enhancing generation quality, reducing hallucinations, and improving alignment between user queries and retrieved content.
+Our results show consistent improvements over strong baselines across diverse domains and context lengths.
+Future work can extend CausalRAG in several directions.
+An important future direction is to ideally evaluate scalability under long-context scenarios—requiring the construction and processing of numerous high-dimensional graphs, each derived from large-scale documents spanning millions of tokens.
+However, current practices often merge documents into a single graph or build smaller graphs for expert-verifiable segments, limiting real-world scalability.
+In addition, as LLMs evolve to support larger input sizes, developing scalable, graph-based causal retrieval methods for long-context reasoning should be an important next step.
 Limitations
@@
 Second, identifying causal paths during inference also requires additional LLM calls, introducing extra computational costs that may limit the performance in real-world deployments.
+Acknowledgement
+We thank the reviewers for their valuable feedback on this work. This research was supported in part by NSF Award (No. 2117439).
 Appendix
+Appendix A LLM Prompts Used
+For transparency we reproduce the exact templates used in our pipeline (see Figure 7).
+A.1 Causal–discovery prompt
+---Role--- You are a smart assistant …
+--- Goal --- Write a structured, professional causality analysis report …
+---Network Data--- {graph_data}
+--- Report Format --- … (see listing in the figure)
+A.2 Causal–summary prompt
+---Role--- You are a helpful assistant …
+---Goal--- Generate a response … merge the cleaned information …
+---Causal Summary--- {causal_summary}
+---Target Response Length and Format--- {response_type}
+---User Query--- {query}
+The prompts follow three practical heuristics: (i) role conditioning to steer style; (ii) explicit section headings to simplify post-processing; (iii) a strict token budget to keep latency manageable.
 Figure 7: LLM prompts for causal discovery and causal summary
-Appendix A LLM Prompts Used
-The full prompts used in the CausalRAG framework for causal discovery and causal summarization (Figure 7), as illustrated in Figure 2. The prompts are adapted from Edge et al. (2024).
-Appendix B OpenAlex Dataset Example
-This is an exmaple (see Table 2) in the dataset Priem et al. (2022b) along with generated questions.
+Appendix B OpenAlex Dataset Construction
+We sample journal and conference papers with full text from the OpenAlex API and stratify by top-level field to avoid domain bias (see Table 1).
+For each document we ask an LLM to draft five answerable questions, inspired by the procedure in Edge et al.
+(2024).
+Table 2 in this appendix shows one complete example of the resulting {document, questions} record.
 Dataset OpenAlex Priem et al. (2022b)
@@
 Table 2: Example document metadata, full-text excerpt, and evaluation questions used in our study.
+Appendix C Implementation Details I – System Configuration
+  • Base LLM. All retrieval-time and evaluation calls use gpt-4o-mini via the standard OpenAI API.
+  • Embeddings. Chunks and queries are encoded with sentence-transformers all-MiniLM-L6-v2 (384 dimension, cosine similarity).
+  • Vector store. Embeddings are stored in a FAISS index; retrieval depth k=3k\!=\!3 throughout.
+  • Orchestration. Pipeline components are wired together with LangChain v0.2; we use its async OpenAI client to parallelise requests.
+Appendix D Implementation Details II – Dataset and Evaluation
+Question generation.
+For every document we instruct the LLM to produce five grounded questions whose answers appear verbatim in the text.
+Constraining the prompts this way keeps the QA task focussed on retrieval quality rather than open-ended speculation.
+Graph creation.
+To assess reasoning on structured knowledge, we constructed individual graphs for each document.
+Although the per-document build is computationally expensive, it largely eliminates information bleed between external knowledge and gives every system the same, self-contained knowledge base.
+Evaluation workload.
+The experiment therefore consists of disjoint {document, question, graph} triples.
+For each triple we evaluate five variants running one retrieval step, one generation step, the $n$ questions, and three automatic metrics (faithfulness, context-precision, context-recall) implemented with the Ragas framework.
+This uniform process enables a fair, like-for-like comparison of different RAGs’ reasoning ability.